OpenStudy (anonymous):

cos x cos y = 1/2(sin(x + y)+ sin(x-y)) is this true or false I went with false I was looking at the product sum

OpenStudy (anonymous):

Which do you think it is?

OpenStudy (anonymous):

\[\begin{cases}\cos (x+y)=\cos x\cos y-\sin x\sin y\\ \cos(x-y)=\cos x\cos y+\sin x\sin y\end{cases}\] Adding the equations gives \[\cos(x+y)+\cos(x-y)=2\cos x\cos y\]